OpenStudy (anonymous):

solve plzz e^x=-3 in R

Parth (parthkohli):

Use logarithms.

OpenStudy (anonymous):

yes but ln(-3) ???!!!!!

Parth (parthkohli):

\[e^x = y \iff \ln y = x \]

OpenStudy (anonymous):

this is my problem ^_^

OpenStudy (anonymous):

 e^x only yields positive numbers

Parth (parthkohli):

Yes, that's right. And so there's no solution in \(\mathbb{R}\)

Parth (parthkohli):

In fact, any exponential function does.

OpenStudy (anonymous):

thanx for the help that's what i was looking for

Parth (parthkohli):

Yay :-)